Project A has cash flows of −$92,000, $49,400, $27,200, $24,500, and $30,690 for Years 0 to 4, respectively. Project B has an initial cost of $50,000 and an annual cash inflow of $20,500 for four years. The required rate of return is 18 percent. These are mutually exclusive projects. Which project(s) should be accepted or rejected?
Answer #1

The project with higher NPV should be selected

NPV = present value of cash inflows - present value of cash outflows

Project A = -92000 + 49400*PVF(18%, 1 year) + 27200*PVF(18%, 2 years) + 24500*PVF(18%, 3 years) + 30,690*PVF(18%, 4 years)

= -92000 + 49400*0.847 + 27200*0.718 + 24500*0.609 + 30690*0.516

= $127.94

Project B = -50,000 + 20,500*2.690

= $5,145

Hence, Project B should be accepted and A should be rejected since the projects are mutually exclusive and only one can be selected